# references.bib

### v2

@@ Abstract
Solving partial differential equations with the finite element method leads to large linear systems of equations that must be solved. When these systems have a natural block structure due to multiple field variables, using iterative solvers with carefully designed preconditioning strategies that exploit the underlying physical structure becomes necessary for an efficient and scalable solution process. FEniCSx Preconditioning Tools (FEniCSx-pctools) is a software package that eases the specification of PETSc (Portable, Extensible Toolkit for Scientific Computation) block preconditioning strategies on linear systems assembled using the DOLFINx finite element solver of the FEniCS Project. The package automatically attaches all necessary metadata so that preconditioning strategies can be applied via PETSc’s standard options-based configuration system. The documented examples include a simple mixed Poisson system and more complex pressure convection-diffusion approach to preconditioning the Navier–Stokes equations. We show weak parallel scaling on a fully coupled temperature-Navier–Stokes system up to $8192$ MPI (Message Passing Interface) processes, demonstrating the applicability of the approach to large-scale problems.
@@ Introduction
Solving linear systems is a core step in the numerical solution of partial differential equations (PDE). To solve large linear systems efficiently on high-performance computers, iterative solvers are essential. However, iterative solvers require effective preconditioning to be viable; without this, convergence may be unacceptably slow, or fail [allaire_numerical_2008]. When modelling complex physical phenomena involving multiple coupled fields, such as incompressible fluid flow, mixed formulations of elasticity, and electromagnetism, the linear systems have a block structure where the different variables have natural couplings. It is now well established that the iterative solution of block linear systems requires block preconditioning strategies that exploit the structure of the coupled problems [benzi_numerical_2005, elman_finite_2014].
The purpose of this software metadata paper is to describe FEniCSx Preconditioning Tools (henceforth FEniCSx-pctools), an add-on for DOLFINx [baratta_dolfinx_2023] and PETSc (Portable, Extensible Toolkit for Scientific Computation) [petsc-user-ref] that eases the specification of complex PETSc-based block-structured preconditioners and linear solvers. FEniCSx-pctools requires both DOLFINx and PETSc/petsc4py and was produced as a supporting tool in a larger research project focused on the topology optimisation of fluidic devices.
The main contribution of FEniCSx-pctools is to bridge the gap between DOLFINx and PETSc by offering a set of algorithms that can analyse the high-level Unified Form Language (UFL) [alnaes_unified_2014] representation of a block-structured finite element problem and subsequently attach the necessary PETSc metadata to the existing DOLFINx assembled linear algebra objects. With this metadata, advanced block preconditioning strategies can be specified straightforwardly using PETSc’s standard options-based configuration system [brown_composable_2012].
We include three fully documented demos. The first shows a Schur complement preconditioner for the mixed Poisson problem, based on a design proposed in [rusten_interior_1996]. The second sets up a Schur complement preconditioner of the velocity-pressure Navier-Stokes equations using the pressure-convection-diffusion (PCD) approach proposed in [elman_finite_2014]. In the final demo, an elementary system of algebraic equations is used to illustrate the ability to change the solver configuration at runtime independently of the model formulation.
@@
We focus on software that provides a high-level interface to block preconditioning strategies in lower-level sparse linear algebra libraries like Trilinos [trilinos-website, baker_tpetra_2012] and PETSc [petsc-user-ref].
@@
    The Firedrake Project [rathgeber_firedrake_2016] builds directly on top of the PETSc DM package [lange_efficient_2016] allowing for the straightforward specification of block-structured algebraic systems and composable physics-based preconditioners.
@@
    CBC.Block [mardal_block_2012] provides block preconditioning tools within the legacy DOLFIN library [logg_dolfin_2010, alnaes_fenics_2015] using the Trilinos linear algebra backend. A particularly strong aspect of CBC.Block is its domain-specific algebraic language for specifying block linear algebra preconditioners. CBC.Block is one of the core components of HAZniCS [budisa_haznics_2023], a software toolbox for solving interface-coupled multiphysics problems.
@@
    FENaPack [fenapackpaper] is a preconditioning package for the legacy DOLFIN library using the PETSc linear algebra backend. A particular focus is on implementing the PCD approach for preconditioning the Navier-Stokes equations.
@@
    PFIBS [chang_pfibs_2022], like FENaPack, is a parallel preconditioning package for the legacy DOLFIN library using the PETSc linear algebra backend. It contains a class BlockProblem that can split a monolithic PETSc matrix into blocks. PFIBS also contains an example of PCD preconditioning of the Navier-Stokes equations.
The first-mentioned finite element solver, Firedrake [rathgeber_firedrake_2016], offers block preconditioning as standard. The remaining packages are extensions to the legacy DOLFIN finite element solver that ease block preconditioning in Trilinos or PETSc. FEniCSx-pctools is similar in scope to these remaining packages, in that it extends the basic capabilities of the new DOLFINx finite element solver [baratta_dolfinx_2023], which is linear algebra backend agnostic and does not use PETSc’s DM functionality like Firedrake, to support PETSc-based block preconditioning. Additionally, like FENaPack and PFIBS, FEniCSx-pctools contains an implementation of PCD preconditioning which is non-trivial to implement for users.
@@
In this section we briefly outline the structure of FEniCSx-pctools library before illustrating its use via the example of preconditioning a mixed Poisson problem. We assume basic knowledge of the mathematics of the finite element method, the DOLFINx finite element solver and PETSc, see e.g. [fenicsx_tutorial_2025].
FEniCSx-pctools is implemented as a Python package fenicsx-pctools with two subpackages named mat (matrix, abbreviated as in PETSc to mat) and pc (preconditioners, abbreviated as in PETSc to pc), and a module utils containing utilities which we will not discuss here. The package is fully documented and is type-hinted/checked.
@@
The pc subpackage primarily contains a class WrappedPC, a PETSc Python-context PC that allows for preconditioners to work with the PETSc Python-context Mat created by the factory functions in the aforementioned mat subpackage. The class WrappedPC is not used directly; instead, users tell PETSc to use it as a Python-context preconditioner and combine it with a standard PETSc PC. More specifically, the Python-context preconditioner sets the standard preconditioner as its own attribute and uses it to interact with the underlying block matrix. This way it can wrap any type of preconditioner (hence the generic name), but it is designed to extend the configurability of the “fieldsplit” preconditioner. Additionally, pc contains implementations of two PCD preconditioning strategies PCDPC_vX and PCDPC_vY. All of the prior mentioned preconditioner classes inherit from a shared base class PCBase which can be used by advanced users to create their own preconditioners.
We now move onto an example showing the use of create_splittable_matrix_block and WrappedPC. A classic example of a block-structured linear system arises from the finite element discretisation of the mixed Poisson problem. Preconditioning this problem is a foundational example and so is widely used to show the basic aspects of preconditioning in software frameworks. In weak form we seek a vector-valued flux $q_{h}\in Q_{h}$ and a scalar-valued pressure $p_{h}\in P_{h}$, such that
$\left(q_{h},\tilde{q}\right)+\left(p_{h},\mathrm{div}(\tilde{q})\right)$ $=0\quad$ $\forall\tilde{q}\in Q_{h},$ (1a)
$\left(\mathrm{div}(q_{h}),\tilde{p}\right)$ $=\left(-f,\tilde{p}\right)\quad$ $\forall\tilde{p}\in P_{h},$ (1b)
@@
where $A$ is a square matrix arising from the bilinear form $(q_{h},\tilde{q})$, $B$ and $B^{T}$ are non-square matrices arising from the bilinear forms $(\mathrm{div}(q_{h}),\tilde{p})$ and $(p_{h},\mathrm{div}(\tilde{q}))$ respectively, $O$ is a square matrix of zeros, 00 is a vector of zeros and $g$ is a vector arising from the linear form $(-f,\tilde{p})$. Finally, we denote the block matrix on the left of eq. 2 $K$, the block vector of flux and pressure unknowns $x$, and the block vector on the right $b$, i.e. $Kx=b$.
The matrix $K$ and the vector $b$ can be assembled in DOLFINx using the standard code shown in fig. 1.
@@
8from dolfinx.fem.petsc import assemble_matrix, assemble_matrix_block, assemble_vector_block
@@
10from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, avg, div, grad, inner, jump
@@
32L = [inner(fem.Constant(domain, (0.0, 0.0)), q_t) * dx, -inner(f, p_t) * dx]
@@
36K = assemble_matrix_block(a_dolfinx)
@@
39b = assemble_vector_block(L_dolfinx, a_dolfinx)
@@
At this point DOLFINx has assembled the left-hand side matrix of eq. 2 in two steps:
  1. 1.
  2. 2.
@@
The extraction of a submatrix is the key mechanism exploited by PETSc’s fieldsplit-based preconditioners [brown_composable_2012]. The splittable block matrix created with the dedicated factory function, see fig. 2, is equipped with metadata including row and column index sets defining the individual blocks. These can be arbitrarily combined (concatenated) and used to define a fieldsplit group to be extracted. The grouping is not limited to neighbouring blocks, which is the case when using PETSc’s “nest” matrices directly with the “fieldsplit” preconditioner.
@@
52from dolfinx.fem.petsc import assemble_matrix, assemble_matrix_block, assemble_vector_block
@@
54from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, avg, div, grad, inner, jump
@@
for $P_{\mathrm{upper}}^{-1}$ in eq. 5, where the tilde $(\tilde{\cdot})$ denotes an approximate (inexact) inverse.
@@
Figure 4: Continuation of fig. 3. We tell PETSc to approximate $\tilde{A}^{-1}$ using one application of block Jacobi preconditioning with the matrix $A$, the already-assembled upper-left block of $K$.
For $\tilde{S}^{-1}$ we take a single application of algebraic multigrid preconditioned discontinuous Galerkin approximation of the Laplacian on the finite element pressure space $P_{h}$. This can be specified as in fig. 5. This choice is justified by the spectral equivalence between the Schur complement and the Laplacian, for further details see e.g. [benzi_numerical_2005].
@@
77from dolfinx.fem.petsc import assemble_matrix, assemble_matrix_block, assemble_vector_block
@@
79from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, avg, div, grad, inner, jump
@@
101L = [inner(fem.Constant(domain, (0.0, 0.0)), q_t) * dx, -inner(f, p_t) * dx]
@@
105K = assemble_matrix_block(a_dolfinx)
@@
108b = assemble_vector_block(L_dolfinx, a_dolfinx)
@@
112K_splittable.setOptionsPrefix("mp_")
113
114solver = PETSc.KSP().create(MPI.COMM_WORLD)
@@
116from mpi4py import MPI
117from petsc4py import PETSc
118
119import numpy as np
@@
We remark that the same configuration of the linear solver can be achieved using the alternative DOLFINx’s assembly routines that produce the PETSc block-structured matrix of type “nest” [brown_composable_2012], and without the use of the utilities provided by FEniCSx-pctools. The main advantage offered by FEniCSx-pctools is the possibility to change the preconditioner setup at runtime without the need to modify the model specification (typically, the arrangement of finite element function spaces that determines the block layout of the system matrix).
@@
We perform weak scalability tests up to $8192$ MPI (Message Passing Interface) processes on a three-field (temperature, velocity and pressure) problem, describing a steady-state incompressible flow with thermal convection, preconditioned with algebraic multigrid for the temperature block and PCD approach for the velocity-pressure Navier-Stokes blocks [elman_finite_2014]. The PDE and problem data (domain, boundary conditions etc.) are exactly the same as those described as the Rayleigh-Bénard problem in [kirby_solver_2018] and solved with Firedrake [rathgeber_firedrake_2016], so for brevity’s sake we do not repeat the details. The solution is described and visualised in fig. 7. Compared with [kirby_solver_2018] we use a slightly different design for the PCD component of our preconditioner following [blechta_towards_2019]. This results in fewer outer Newton-Raphson iterations and fewer inner Krylov solver iterations than the design proposed in [kirby_solver_2018].
@@
Figure 7: Visualisation of the solution to the Rayleigh-Bénard problem [kirby_solver_2018]. The fluid velocity is shown as streamlines coloured by the velocity magnitude, the fluid pressure as an isocontour field, and the fluid temperature on the back surface only. Fluid is held at a higher temperature on the left and a lower temperature on the right of the cube. Less dense fluid rises on the left, and denser fluid sinks on the right (the setup takes the gravitational acceleration vector pointing up). At steady-state this forms a distinctive circulation pattern which is clearly shown in the fluid velocity streamlines.
@@
The weak scalability tests were performed on the University of Luxembourg Aion cluster [varrette_management_2022]. The Aion cluster consists of 354 nodes each with two 64 core AMD Epyc ROME 7H12 processors attached to 256 GB RAM. The nodes are connected with an Infiniband HDR network in a ‘fat-tree’ topology. We built DOLFINx 0.7.0 against PETSc 3.20.0, OpenMPI 4.0.5 and Python 3.8.6 using the GCC 10.2.0 compiler suite with -march=znver2 -O3 optimisation flags. All experiments were performed at 50% core utilisation per node, i.e. 64 cores per node, as these low-order finite element problems are typically memory-bandwidth constrained. The node allocation was taken in exclusive mode, i.e. with no other jobs running on the same node.
@@
In the weak scaling test we solve the problem on 1 node and double the number of nodes until we reach 128 nodes ($8192$ MPI processes). We simultaneously increase the problem size so that the number of degrees of freedom (DOF) per MPI rank remains fixed at around 100 000100000100\,000100 000. A breakdown of performance metrics for the weak scaling study is shown in table 1. The last column shows the total wall time to execute the SNES.solve method, which includes the DOLFINx assembly of the necessary linear operators and their PETSc preconditioning and solution. In summary, the time for solution stays roughly constant, demonstrating the excellent combined weak scaling of DOLFINx, PETSc and the coupling implemented by FEniCSx-pctools. The code to execute these experiments and our raw timing data is available in the supplementary material in the examples/rayleigh-benard-convection directory.
@@
(×$\times 10^{6}$) processes iterations iterations iterations iterations SNES solve (s)
@@
Table 1: Performance metrics for the Rayleigh-Bénard problem [kirby_solver_2018] with customised PCD-AMG preconditioning. Weak scaling at 100k (DOF) per process. Aion Cluster, 50% utilisation. The number in brackets is the average iterations per outer linear solve. Wall time for SNES solve is the time taken to execute SNES.solve, which includes PETSc linear algebra operations, DOLFINx assembly operations and the coupling implemented by FEniCSx-pctools.
@@
FEniCSx-pctools is written in Python and is compatible with the CPython intepreter version 3.9 and above.
@@
The design of parameter and discretisation robust block preconditioning strategies is an active research topic in numerical analysis and computational sciences. We can point to recent developments in designs for the Navier–Stokes [farrell_reynolds-robust_2021] equations, poroelasticity equations [chen_robust_2020], magnetohydrodynamic equations [laakmann_augmented_2022] and multiphysics interface problems [budisa_rational_2023]. Together, FEniCSx-pctools, DOLFINx and PETSc support the straightforward expression and testing of these preconditioning strategies in code, and therefore are useful for researchers who wish to quickly verify the performance of their preconditioning designs. In addition, block preconditioning strategies are an important tool for solving large real-world problems in computational sciences and engineering.
@@
  • Version published: Version 5 created from git tag v0.9.1
  • Date published: 08/07/2025 (ongoing)

### v3

# FEniCSx-pctools: Tools for PETSc block linear algebra preconditioning in FEniCSx

@@ Abstract
Solving partial differential equations with the finite element method leads to large linear systems of equations that must be solved. When these systems have a natural block structure due to multiple field variables, using iterative solvers with carefully designed preconditioning strategies that exploit the underlying physical structure becomes necessary for an efficient and scalable solution process. FEniCSx Preconditioning Tools (FEniCSx-pctools) is a software package that eases the specification of PETSc (Portable, Extensible Toolkit for Scientific Computation) block preconditioning strategies on linear systems assembled using the DOLFINx finite element solver of the FEniCS Project. The package automatically attaches all necessary metadata so that preconditioning strategies can be applied via PETSc’s standard options database to monolithic and block assembled systems. The documented examples include a simple mixed Poisson system and more complex pressure convection-diffusion approach to preconditioning the Navier-Stokes equations. We show weak parallel scaling on a fully coupled temperature-Navier-Stokes system up to $8192$ MPI (Message Passing Interface) processes, demonstrating the applicability of the approach to large-scale problems.
@@ Introduction
Solving linear systems is a core step in the numerical solution of partial differential equations (PDE). To solve large linear systems efficiently on high-performance computers, iterative solvers are essential. However, iterative solvers require effective preconditioning to be viable; without this, convergence may be unacceptably slow, or fail [1]. When modelling complex physical phenomena involving multiple coupled fields, such as incompressible fluid flow, mixed formulations of elasticity, and electromagnetism, the linear systems have a block structure where the different variables have natural couplings. It is now well established that the iterative solution of block linear systems requires block preconditioning strategies that exploit the structure of the coupled problems [7, 16].
The purpose of this software metadata paper is to describe FEniCSx Preconditioning Tools (henceforth FEniCSx-pctools), an add-on for DOLFINx [6] and PETSc (Portable, Extensible Toolkit for Scientific Computation) [5] that eases the specification of complex PETSc-based block-structured preconditioners and linear solvers. FEniCSx-pctools requires both DOLFINx and PETSc/petsc4py and was produced as a supporting tool in a larger research project focused on the topology optimisation of fluidic devices.
The main contribution of FEniCSx-pctools is to bridge the gap between DOLFINx and PETSc by offering a set of algorithms that can analyse the high-level Unified Form Language (UFL) [3] representation of a block-structured finite element problem and subsequently attach the necessary PETSc metadata to the existing DOLFINx assembled monolithic and block assembled linear algebra objects. With this metadata, advanced block preconditioning strategies can be specified straightforwardly using PETSc’s standard options-based configuration system [10].
We include three fully documented demos. The first shows a Schur complement preconditioner for the mixed Poisson problem, based on a design proposed in [24]. The second sets up a Schur complement preconditioner of the velocity-pressure Navier-Stokes equations using the pressure-convection-diffusion (PCD) approach proposed in [16]. In the final demo, an elementary system of algebraic equations is used to illustrate the ability to change the solver configuration at runtime independently of the model formulation.
@@
We focus on software that provides a high-level interface to block preconditioning strategies in lower-level sparse linear algebra libraries like Trilinos [25, 4] and PETSc [5].
@@
    The Firedrake Project [23] builds directly on top of the PETSc DM package [20] allowing for the straightforward specification of block-structured algebraic systems and composable physics-based preconditioners.
@@
    CBC.Block [22] provides block preconditioning tools within the legacy DOLFIN library [21, 2] using the Trilinos linear algebra backend. A particularly strong aspect of CBC.Block is its domain-specific algebraic language for specifying block linear algebra preconditioners. CBC.Block is one of the core components of HAZniCS [11], a software toolbox for solving interface-coupled multiphysics problems.
@@
    FENaPack [9] is a preconditioning package for the legacy DOLFIN library using the PETSc linear algebra backend. A particular focus is on implementing the PCD approach for preconditioning the Navier-Stokes equations.
@@
    PFIBS [13], like FENaPack, is a parallel preconditioning package for the legacy DOLFIN library using the PETSc linear algebra backend. It contains a class BlockProblem that can split a monolithic PETSc matrix into blocks. PFIBS also contains an example of PCD preconditioning of the Navier-Stokes equations.
The first-mentioned finite element solver, Firedrake [23], offers block preconditioning as standard. The remaining packages are extensions to the legacy DOLFIN finite element solver that ease block preconditioning in Trilinos or PETSc. FEniCSx-pctools is similar in scope to these remaining packages, in that it extends the basic capabilities of the new DOLFINx finite element solver [6], which is linear algebra backend agnostic and does not use PETSc’s DM functionality like Firedrake, to support PETSc-based block preconditioning. Additionally, like FENaPack and PFIBS, FEniCSx-pctools contains an implementation of PCD preconditioning which is non-trivial to implement for users.
@@
In this section we briefly outline the structure of FEniCSx-pctools library before illustrating its use via the example of preconditioning a mixed Poisson problem. We assume basic knowledge of the mathematics of the finite element method, the DOLFINx finite element solver and PETSc, see e.g. [15].
Before continuing we briefly explain the three main types of matrix assembly routines that DOLFINx offers in the current development version 0.10.0.dev0 and whether they allow for block preconditioning.
The first type, monolithic, is created when assembling via assemble_matrix(..., kind= None) a finite element form defined on a single dolfinx.fem.FunctionSpace containing a single basix.ufl.FiniteElement or a single basix.ufl.MixedElement containing multiple basix.ufl.FiniteElement objects (one for each variable). In the latter case, the degree-of-freedom maps related to each variable are interleaved into a single global degree-of-freedom map, resulting in strong data locality, but the assembled matrix loses the ‘natural’ block structure and block preconditioning is not available.
The second type, nest, is created when assembling a finite element form defined on a sequence of dolfinx.fem.FunctionSpace objects using the assemble_matrix(..., kind="nest") call. Each interaction between a pair of variables (including self-interactions) is assembled into its own matrix and the global system is represented as a PETSc MATNEST matrix composed of these submatrices. The advantage of the nest type is that block preconditioning is supported natively, but in the case the user wants to use a direct solver, only MUMPS can be used.
The third type, block, is created when assembling a finite element form defined on a sequence of dolfinx.fem.FunctionSpace objects using the assemble_matrix(..., kind="mpi") call. In this case, the degree-of-freedom maps for each variable occupy a contiguous block in a global degree-of-freedom map. The forms are assembled into a standard PETSc MATMPIAIJ matrix and the block structure is retained. An advantage of the block type is that any direct solver can be used for solution, not just MUMPS, as in nest. However, as the degree-of-freedom map has lost explicit information about the variables, block preconditioning is not available.
The key contribution of FEniCSx-pctools is to enable block preconditioning for monolithic and block assembly.
FEniCSx-pctools is implemented as a Python package fenicsx-pctools with two subpackages named mat (matrix, abbreviated as in PETSc to mat) and pc (preconditioners, abbreviated as in PETSc to pc). The package is fully documented and is type-hinted/checked.
@@
The pc subpackage primarily contains a class WrappedPC, a PETSc Python-context PC that allows for preconditioners to work with the PETSc Python-context Mat created by the factory functions in the aforementioned mat subpackage. The class WrappedPC is not used directly; instead, users tell PETSc to use it as a Python-context preconditioner and combine it with a standard PETSc PC. More specifically, the Python-context preconditioner sets the standard preconditioner as its own attribute and uses it to interact with the underlying block matrix. This way it can wrap any type of preconditioner (hence the generic name), but it is designed to extend the configurability of the fieldsplit preconditioner. Additionally, pc contains implementations of two PCD preconditioning strategies PCDPC_vX and PCDPC_vY. All of the prior mentioned preconditioner classes inherit from a shared base class PCBase which can be used by advanced users to create their own preconditioners.
The diagram in fig. 1 illustrates the role of FEniCSx-pctools in the FEniCS ecosystem, and shows how to set up a fieldsplit preconditioner to allow block preconditioning of a block matrix.
Figure 1: Diagram summarising the workflow to implement a custom fieldsplit preconditioner using the FEniCSx-pctools library. (Top left block, UFL) The symbolic UFL bilinear form defining a block operator is first defined as a 2D list of UFL forms a_ufl. (Top right block, DOLFINx) The list of UFL forms is just-in-time compiled to a list of compiled finite element forms a_dolfinx, which is then assembled into a block matrix K in PETSc’s “mpiaij” format (bottom right block, PETSc/petsc4py). We now enter the domain of FEniCSx-pctools, bottom left block. Instead of directly associating the matrix K with the linear solver, we first pass the original UFL form a_ufl and K to create_splittable_matrix_block to create a splittable matrix K_splittable. This is a necessary step to allow block extraction based on arbitrarily combined index sets provided by the fieldsplit preconditioner fs_wrapped. The extraction itself is implemented in the createSubMatrix method of the splittable matrix’ Python-context K_ctx, while the index sets are configurable via setUp method of the preconditioner’s Python-context fs_ctx. (Bottom right block, PETSc/petsc4py) The methods in fs_ctx and K_ctx are transparently executed by the linear solver upon calling its solve method with the right-hand side vector b and a vector x to store the solution.
We now move onto an example for DOLFINx version 0.10.0.dev0 showing the use of create_splittable_matrix_block and WrappedPC . A classic example of a block-structured linear system arises from the finite element discretisation of the mixed Poisson problem. Preconditioning this problem is a foundational example and so is widely used to show the basic aspects of preconditioning in software frameworks. In weak form we seek a vector-valued flux $q_{h}\in Q_{h}$ and a scalar-valued pressure $p_{h}\in P_{h}$, such that
@@
[…]
where $\left(\cdot,\cdot\right)$ denotes the usual $L^{2}$ inner product on the finite element mesh and $f$ a known forcing term. In discrete block form this can be written as a saddle point linear system with unknown vectors of finite element coefficients $q$ and $p$
@@
[…]
The matrix $K$ and the vector $b$ can be assembled in DOLFINx using the standard code shown in fig. 2. At this point DOLFINx has assembled the left-hand side matrix of eq. 2 in two steps:
  1. 1.
  2. 2.
@@
8from dolfinx.fem.petsc import assemble_matrix, assemble_vector
@@
10from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, ZeroBaseForm, avg, div, ds, dS, grad, inner, jump
@@
32L = [ZeroBaseForm((q_t,)), -inner(f, p_t) * dx]
@@
36K = assemble_matrix(a_dolfinx, kind="mpi")
@@
39b = assemble_vector(L_dolfinx, kind="mpi")
@@
The extraction of a submatrix is the key mechanism exploited by PETSc’s fieldsplit-based preconditioners [10]. The splittable block matrix created with the dedicated factory function, see fig. 3, is equipped with metadata including row and column index sets defining the individual blocks. These can be arbitrarily combined (concatenated) and used to define a fieldsplit group to be extracted. The grouping is not limited to neighbouring blocks, which is the case when using PETSc MATNEST matrices directly with the fieldsplit preconditioner.
@@
52from dolfinx.fem.petsc import assemble_matrix, assemble_vector
@@
54from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, ZeroBaseForm, avg, div, ds, dS, grad, inner, jump
@@
for $P_{\mathrm{upper}}^{-1}$ in eq. 5, where the tilde (~)($\tilde{\cdot}$) denotes an approximate (inexact) inverse.
@@
Figure 5: Continuation of fig. 4. We tell PETSc to approximate $\tilde{A}^{-1}$ using one application of block Jacobi preconditioning with the matrix AA, the already-assembled upper-left block of $K$.
@@
77from dolfinx.fem.petsc import assemble_matrix, assemble_vector
@@
79from ufl import CellDiameter, FacetNormal, Measure, TestFunction, TrialFunction, ZeroBaseForm, avg, div, ds, dS, grad, inner, jump
@@
101L = [ZeroBaseForm((q_t,)), -inner(f, p_t) * dx]
@@
105K = assemble_matrix(a_dolfinx, kind="mpi")
@@
108b = assemble_vector(L_dolfinx, kind="mpi")
@@
113from mpi4py import MPI
114from petsc4py import PETSc
115
116import numpy as np
@@
For $\tilde{S}^{-1}$ we take a single application of algebraic multigrid preconditioned discontinuous Galerkin approximation of the Laplacian on the finite element pressure space $P_{h}$. This can be specified as in fig. 6. This choice is justified by the spectral equivalence between the Schur complement and the Laplacian, for further details see e.g. [7]. In the final step, we complete the solver setup and solve the problem using the sequence of commands shown in fig. 7.
We remark that the same configuration of the linear solver can be achieved using the alternative DOLFINx’s assembly routines that produce the PETSc block-structured matrix of type MATNEST [10], and without the use of the utilities provided by FEniCSx-pctools. The main advantage offered by FEniCSx-pctools is the possibility to change the preconditioner setup at runtime without the need to modify the model specification (typically, the arrangement of finite element function spaces that determines the block layout of the system matrix).
@@
We perform weak scalability tests up to $8192$ MPI (Message Passing Interface) processes on a three-field (temperature, velocity and pressure) problem, describing a steady-state incompressible flow with thermal convection, preconditioned with algebraic multigrid for the temperature block and PCD approach for the velocity-pressure Navier-Stokes blocks [16]. The PDE and problem data (domain, boundary conditions etc.) are exactly the same as those described as the Rayleigh-Bénard problem in [18] and solved with Firedrake [23], so for brevity’s sake we do not repeat the details. The solution is described and visualised in fig. 8. Compared with [18] we use a slightly different design for the PCD component of our preconditioner following [8]. This results in fewer outer Newton-Raphson iterations and fewer inner Krylov solver iterations than the design proposed in [18].
@@
Figure 8: Visualisation of the solution to the Rayleigh-Bénard problem [18]. The fluid velocity is shown as streamlines coloured by the velocity magnitude, the fluid pressure as an isocontour field, and the fluid temperature on the back surface only. Fluid is held at a higher temperature on the left and a lower temperature on the right of the cube. Less dense fluid rises on the left, and denser fluid sinks on the right (the setup takes the gravitational acceleration vector pointing up). At steady-state this forms a distinctive circulation pattern which is clearly shown in the fluid velocity streamlines.
@@
The weak scalability tests were performed on the University of Luxembourg Aion cluster [26]. The Aion cluster consists of 354 nodes each with two 64 core AMD Epyc ROME 7H12 processors attached to 256 GB RAM. The nodes are connected with an Infiniband HDR network in a ‘fat-tree’ topology. We built DOLFINx 0.7.0 against PETSc 3.20.0, OpenMPI 4.0.5 and Python 3.8.6 using the GCC 10.2.0 compiler suite with -march=znver2 -O3 optimisation flags. All experiments were performed at 50% core utilisation per node, i.e. 64 cores per node, as these low-order finite element problems are typically memory-bandwidth constrained. The node allocation was taken in exclusive mode, i.e. with no other jobs running on the same node.
@@
In the weak scaling test we solve the problem on 1 node and double the number of nodes until we reach 128 nodes ($8192$ MPI processes). We simultaneously increase the problem size so that the number of degrees of freedom (DOF) per MPI rank remains fixed at around 100 000100\,000. A breakdown of performance metrics for the weak scaling study is shown in table 1. The last column shows the total wall time to execute the SNES.solve method, which includes the DOLFINx assembly of the necessary linear operators and their PETSc preconditioning and solution. In summary, the time for solution stays roughly constant, demonstrating the excellent combined weak scaling of DOLFINx, PETSc and the coupling implemented by FEniCSx-pctools. The code to execute these experiments and our raw timing data is available in the supplementary material in the examples/rayleigh-benard-convection directory.
@@
($\times 10^{6}$) processes iterations iterations iterations iterations SNES solve (s)
@@
Table 1: Performance metrics for the Rayleigh-Bénard problem [18] with customised PCD-AMG preconditioning. Weak scaling at 100k (DOF) per process. Aion Cluster, 50% utilisation. The number in brackets is the average iterations per outer linear solve. Wall time for SNES solve is the time taken to execute SNES.solve, which includes PETSc linear algebra operations, DOLFINx assembly operations and the coupling implemented by FEniCSx-pctools.
@@
FEniCSx-pctools is written in Python and is compatible with the CPython intepreter version 3.10 and above.
@@
The design of parameter and discretisation robust block preconditioning strategies is an active research topic in numerical analysis and computational sciences. We can point to recent developments in designs for the Navier-Stokes [17] equations, poroelasticity equations [14], magnetohydrodynamic equations [19] and multiphysics interface problems [12]. Together, FEniCSx-pctools, DOLFINx and PETSc support the straightforward expression and testing of these preconditioning strategies in code, and therefore are useful for researchers who wish to quickly verify the performance of their preconditioning designs. In addition, block preconditioning strategies are an important tool for solving large real-world problems in computational sciences and engineering.
@@
  • Version published: Version 6 created from git tag v0.10.0.dev0-paper
  • Date published: 04/09/2025 (ongoing)